OpenStudy (anonymous):

An apple ipod is 299 an its marked up by 40% whats the final coast

OpenStudy (perl):

299 ( 1 + .4 )

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[299 \times .6 = x\] \[179.4 = x\] x = the sale price. Hope this helps :)

OpenStudy (anonymous):

418.60

OpenStudy (anonymous):

299 * .4 = 119.6 + 299 = 418.60

OpenStudy (anonymous):

Thank you guys so much. :-):-)

OpenStudy (dwelsher):

299*.4=119.6+299=418.6